# Paper Quality Assessment based on Individual Wisdom Metrics from Open Peer Review

### v1

@@ Acknowledgements
This work has been supported by ONR award N00014-19-1-2584, by NSF-CISE award IIS-2151077 under the Robust Intelligence program, by the ARO-MURI award W911NF-23-1-0277, by the Simons Foundation SCGB program 1181110, and the K. Lisa Yang ICoN Center. All authors would like to thank Laurence Hunt, Mariya Toneva, Steven Scholte, and the organizing committee of the 2023 Conference on Cognitive Computational Neuroscience for their support and feedback throughout the project. A.Z. expresses gratitude to Professors James DiCarlo and Rebecca Saxe from the Department of Brain and Cognitive Sciences at MIT for their feedback, guidance and inspiration in early stages of this project.
@@ Appendix 1. MSD Derivations
MSD(simple $\text{MSD}(\text{simple mean})=\mathbb{E}\left[\left(\frac{1}{n}\sum_{i=1}^{n}\epsilon_{ij}\right)^{2}\right].$
@@
MSD(simple $\text{MSD}(\text{simple mean})=\frac{1}{n^{2}}\sum_{i=1}^{n}\sigma_{i}^{2}$
@@
[…]
@@
[…]
@@
Supplementary Figure S1: Confidence of ratings is inversely correlated to the assigned scores. In the ICLR2023 open peer review data, confidence of assigned ratings was inversely correlated to the assigned scores ($r=-0.115,p<0.001,N=14359$). For every review of every submission in ICLR2023, this plot shows the overall paper score given by the reviewer on the y-axis and the corresponding confidence assigned by the reviewer on the x-axis.
Supplementary Figure S2: Reviewers with intermediate author scores are the best at predicting the community average paper quality score. (A) Distribution of mean standard deviations from community average score across the different reviewer groups, grouped by the authorship score of the reviewer. (B) Reviewers with the author scores in the range between 40th and 60th percentiles have significantly lower mean squared deviation from the community average score than the reviewers from the ends of the spectrum of author scores (0-10th and 90-100th percentiles; $*p<0.05$).
Supplementary Figure S3: Reviewers with intermediate author scores correlate the most with other reviewers within their group. Correlation between pairs of reviewers reviewing the same paper, split by the authorship score of the reviewers. The correlation is the highest for the reviewers with intermediate authorship scores (in the range between 40th and 60th percentiles; $N>90$ pairs, all panels).

### v2

@@ Acknowledgements
This work has been supported by NSF-CISE award IIS-2151077 under the Robust Intelligence program and by the Simons Foundation SCGB program 1181110. All authors would like to thank Laurence Hunt, Mariya Toneva, Steven Scholte, and the organizing committee of the 2023 Conference on Cognitive Computational Neuroscience for their support and feedback throughout the project. A.Z. expresses gratitude to Professors James DiCarlo and Rebecca Saxe from the Department of Brain and Cognitive Sciences at MIT for their feedback, guidance and inspiration in earlier stages of this project.
@@ Appendix 1. MSD Derivations
$\text{MSD}(\text{simple mean})=\mathbb{E}\left[\left(\frac{1}{n}\sum_{i=1}^{n}\epsilon_{ij}\right)^{2}\right].$
@@
$\text{MSD}(\text{simple mean})=\frac{1}{n^{2}}\sum_{i=1}^{n}\sigma_{i}^{2}$
@@
$\text{MSD}(\text{Bayes-optimal})=\mathbb{E}\left[\left(\sum_{i=1}^{n}w_{i}\epsilon_{ij}\right)^{2}\right].$
@@
$\text{MSD}(\text{Bayes-optimal})=\frac{1}{\sum_{i=1}^{n}1/\sigma_{i}^{2}}.$
@@
Supplementary Figure S1: Confidence of ratings is inversely correlated to the assigned scores. In the ICLR2023 open peer review data, confidence of assigned ratings was inversely correlated to the assigned scores (r=0.115,p<0.001,N=14359r=-0.115,p<0.001,N=14359). For every review of every submission in ICLR2023, this plot shows the overall paper score given by the reviewer on the y-axis and the corresponding confidence assigned by the reviewer on the x-axis.
Supplementary Figure S2: Reviewers with intermediate author scores are the best at predicting the community average paper quality score. (A) Distribution of mean standard deviations from community average score across the different reviewer groups, grouped by the authorship score of the reviewer. (B) Reviewers with the author scores in the range between 40th and 60th percentiles have significantly lower mean squared deviation from the community average score than the reviewers from the ends of the spectrum of author scores (0-10th and 90-100th percentiles; p<0.05*p<0.05). XXX indicate which dataset
Supplementary Figure S3: Reviewers with intermediate author scores correlate the most with other reviewers within their group. Correlation between pairs of reviewers reviewing the same paper, split by the authorship score of the reviewers. The correlation is the highest for the reviewers with intermediate authorship scores (in the range between 40th and 60th percentiles; $N>90$ pairs, all panels). XXX indicate which dataset
Supplementary Figure S4: Including a minimum confidence threshold (an upper bound of $\sigma_{\text{total}}$) is necessary, for without it, the correlation between true and estimated paper quality drops, even when using the oracle to estimate the scores.